Could you please answer these questions

1-


Calculate the average density of the following ast


++++++++



2-


Calculate the average density of the following ast

+++++++


3-


Calculate the average density of the following ast

+++++++


4-


Calculate the average density of the following ast



++++++++++++++

5-


Calculate the average density of the following ast


++++++++++

6-


Calculate the average density of the following ast


+++++++++


7-



Calculate the average density of the following ast


+++++++


8-


Calculate the average density of the following ast

++++++++++++++


9-


Calculate the average density of the following ast
Answer #1


1 ) desity of mars = 3.93 g/cm
Answer #2

6 ) R = radius of car   in mile

    number of revolution = 55000 / (2*pi*R )



9 ) a) totoise is X meter from finish line .


then (0.85 -x)/0.14 = ( 0.85- 0.68)/7.7

X = 0.8469 m


totoise is 0.8469 meter from finish line


b ) time = 0.8469/0.14 - 0.68/7.7 = 5.96 s
Answer #4

) desity of mars = 3.93 g/cm
Answer #5

1)D = 3.9242E+12 kg/km^3

2taking fm as femto

a).23*10^18kg/m^3

b)more than mercury

3)9.18nm/s

4)209691.912m^3

5)229253190.01cm^3

229.253m^3

7)

a)-2.5m/sec

b)-2m/sec

c)4 sec

8)19.25 mi/h

9)

a)169.85m

b)1.078sec